# Semiparametric Growth-Curve Modeling in Hierarchical, Longitudinal Studies

### v1

@@ 1 Introduction
An important step in such analysis is the estimation of the underlying growth (decay) curves. Parametric forms like Logistic and Gompertz, as well as nonparametric forms such as smoothing splines, are often used for modeling such monotonic patterns. The usefulness and popularity of parametric approach is clear from the large number of R packages available in the Comprehensive R Archive Network (CRAN), which support parametric growth-curve modeling. A good summary of the state of art in the parametric growth-curve modeling is provided in bioGrow1, which describes the functionality of an R package, biogrowth. In nonparametric growth curve modeling, smoothing splines with and without monotonicity constraints is a popular approach, e.g., see, MonoSpline1, MonoSpline2 and the references therein. In Covid1, the spread of COVID-19 cases is modeled using a state space modeling based approach. They, however, model the growth rate (derivative of the growth curve) rather than the growth curve itself. All these references, with the exception of bioGrow1, are primarily concerned with the modeling of a single growth curve. In hierarchical, longitudinal studies, multiple growth curves are generated as a result of monitoring multiple subjects from a group (or multiple groups). In such cases the interest lies in estimating the group-mean-curves, after accounting for the subject specific differences. For these types of problems the dominant approaches have been based on linear mixed-effects, nonlinear mixed-effects, and functional mixed-effects modeling, e.g., see grimm, blozis. In these references, when linear mixed-effects or functional mixed-effects models are used, the group-mean-curves are often linear or polynomial smoothing splines, which, while flexible, do not use the prior monotonicity information of the growth curves. In nonlinear mixed-effects modeling, it is assumed that the subject curves are generated by parametric growth curve models such as Logistic and some or all of the model parameters, like the growth rate, for each subject are assumed to be random perturbations of their mean population values. However, in this case, the subject curves being fully parametric can be a restrictive assumption.
@@
The group-mean-curve, which is expected to follow a growth curve pattern, is called a functional fixed-effect, the deviation-from-mean-curve is called a functional random-effect, regression-effects can either be fixed or random, and the noise is a zero-mean, Gaussian error term. In this article, the group-mean-curve is modeled semiparametrically according to a simple modification of the state space model based approach described in section 3 of ans1. Our modification makes their approach computationally feasible in much wider settings. Making their elegant growth curve modeling approach easily accessible to the wider applied research community is one of the goals of this article. All the FME models considered in this article are formulated as linear, Gaussian, state space models, SSMs for short. This provides many additional advantages such as modeling flexibility, computational efficiency, and easy interpretability (see Guo1, selu_r:15). The analysis for the illustrations in this article is carried out by using the CSSM procedure in the SAS Viya/Econometrics software, see cssm1. The data and the SAS code used for these illustrations is available on request.
@@ 2 State Space Modeling Framework
SSMs have long been used to model a variety of sequential data, such as time series, panels of time series, and longitudinal data. Many different forms of SSMs are used in practice. In order to handle the formulation of FME models for the longitudinal data, this article uses an SSM form that is similar to the one described in chap. 6, sec. 4 of dk. The notation and additional details are included in the appendix, Appendix A.
@@ 2.1 SSM Formulation of a Parametric Growth Curve
where $\epsilon_{t}$ is a noise sequence of independent, zero-mean, Gaussian variables with variance $\sigma^{2}_{\epsilon}$, and the latent curve, $f(\boldsymbol{\theta},t)=constant+scale\;g(\boldsymbol{\theta},t)$, is a smooth, monotonic function of $t$, which might depend on some parameter vector $\boldsymbol{\theta}$; $constant$ and $scale$ are auxiliary parameters that do not depend on $\boldsymbol{\theta}$. A few polpular choices of $g(\boldsymbol{\theta},t)$ are shown in Table 1; several more examples can be found in zw.
@@
As an illustration, consider the greek_tractors data set considered in bioGrow1. It contains the number of tractors registered in Greece in a 45 year period between 1961 to 2006. They fit the Logistic curve to $y=\log_{10}(\text{number of tractors})$, by nonlinear least squares. After substituting the parameter estimates, their fitted Logistic curve has the following expression:
@@
Note that, in the SSM formulation only $\phi,\rho$ and $\sigma^{2}_{\epsilon}$ are estimated by the maximum likelihood. The other two parameters, $constant$ and $scale$, are obtained as a result of Kalman smoothing. This is because, Kalman smoothing provides the estimates of the two-dimensional latent state $\boldsymbol{\alpha}_{t}$ at all times and $scale$ is the second element of $\boldsymbol{\alpha}_{t}$. Furthermore, since the first element of $\boldsymbol{\alpha}_{t}$ is $f(t)$, once $scale$, $\phi$, and $\rho$ are known, $constant$ can be derived from the estimate of $f(t)$ at some convenient $t$, say at $t=0$. Finally, Figure 1 shows the estimated Logistic fit (which matches the fit in bioGrow1) with point-wise 95% confidence band, which is also provided by the Kalman smoother.
@@ 2.2 SSM Formulation of a Semiparametric Growth Curve
The state space formulations in Equation 2 and Equation 5 are based on the treatment of semiparametric growth-curve modeling described in Section 3 of ans1. There they describe how to obtain $\mathbf{Q}_{t}$ for different choices of $g(\boldsymbol{\theta},t)$ so that the estimated latent curve, $\mu_{t}$, in Equation 4 satisfies a penalized least-squares criteria. However, except for some simple choices of $g(\boldsymbol{\theta},t)$, obtaining a closed form expression for $\mathbf{Q}_{t}$ is complicated. Instead, in this article we suggest always using the following greatly simplified form of $\mathbf{Q}_{t}$:
@@
  • The transition matrix, $\mathbf{T}_{t}^{t+h}$, used in the formulations in Equation 2 and Equation 5, is exactly the same as in Section 3 of ans1.
@@
  • For the exponential case, $g(\boldsymbol{\rho},t)=\exp(-\rho\;t)$, a closed form expression for $\mathbf{Q}_{t}$ is described in ans1 and an equivalent form of this model is discussed in deJong. For this exponential case, we compared the growth/decay curves estimated using the exact form of $\mathbf{Q}_{t}$ versus the simplified form in Equation 6 in many real and simulated data cases. In all cases the estimated curves by the exact and approximate methods matched very well. In particular, for the chlorine content data example discussed in both ans1 and deJong, the estimated curves were essentially identical.
@@ 2.3 FME Model With Growth Curve As a Functional Fixed-Effect
As a quick illustration of FME model based data analysis, we analyze a small portion of the data used in jap. Slightly more elaborate analysis of these data is done in Section 3. Here we consider 12 independent replicate curves generated by monitoring the growth of a bacterium strain in a nutrient rich medium (strain 1, medium LB). The measurements are taken every 30 minutes during a period of 23 hours. Figure 3 shows the growth patterns for replicates 1 and 7. It shows that both the replicates loosely follow the usual sigmoidal shape, however, they do show noticeable differences in their growth patterns.
@@ 3 Illustration: Comparing Bacterial Growth in Different Settings
In jap, correlation between genome reduction and bacterial growth is studied (a link to access their data is also provided in the article). In their study, E. Coli cultures are grown in three settings of growth mediums and at 36 different levels of gene deletion. For each combination of growth medium and gene deletion level, multiple replicates of E. Coli cultures are grown. Their growth is measured at 30 minute or one hour intervals for different lengths of times. The three growth mediums are M63, MAA, and LB, which indicate poor, supplementary, and rich growth conditions, respectively. The 36 gene deletion levels are labeled as strains 1 to 36, which indicate increasing severity of gene deletion. In our illustration we will analyze the growth patterns for strains 1 and 36, grown at all the three mediums, M63, MAA, and LB (that is, total of 6 strain-medium combinations). Except in the case of (strain 36, medium M63) combination, which has 24 replicates, all other combinations have 12 replicates. Even though our methodology can easily handle different number of replicates for different combinations, we excluded the last 12 replicates for the (strain 36, medium M63) combination. Additionally, two minor adjustments to the data were made for easier viewing of the results:
@@ 4 Conclusion
We have described a flexible SSM-based framework for semiparametric modeling of growth curve data that are generated by hierarchical, longitudinal studies. Our modeling approach is based on a simplification of growth curve modeling approach that is described in ans1. We hope our treatment of their approach is more accessible and easier to adopt. The SSM-based framework described here can be easily extended in many directions. For example, the FME model in Equation 7 can contain regression effects, or additional effects that modify the mean curve to account for hormetic effects.
@@ Appendix A SSM Framework and Notation
where the terms $\boldsymbol{\mu}_{t}$, $\boldsymbol{\gamma}_{t},\cdots$ might represent a time-varying mean-level, a seasonal pattern, and so on. Such linear combinations of the state subblocks are called components. When the data from a longitudinal study is assumed to follow an SSM, the data analyis is greately helped by the well-known (diffuse) Kalman filter and (diffuse) Kalman smoother algorithms. Chapters 4, 5, 6, and 7 of dk explain how these algorithms provide the following:

### v2

@@ 1 Introduction
An important step in such analysis is the estimation of the underlying growth (decay) curves. Parametric forms like Logistic and Gompertz, as well as nonparametric forms such as smoothing splines, are often used for modeling such monotonic patterns. The usefulness and popularity of parametric approach is clear from the large number of R packages available in the Comprehensive R Archive Network (CRAN), which support parametric growth-curve modeling. A good summary of the state of art in the parametric growth-curve modeling is provided in Garre et al. [2023], which describes the functionality of an R package, biogrowth. In nonparametric growth curve modeling, smoothing splines with and without monotonicity constraints is a popular approach, e.g., see, Ramsay [1988], Chan et al. [2021] and the references therein. In Harvey and Kattuman [2020], the spread of COVID-19 cases is modeled using a state space modeling based approach. They, however, model the growth rate (derivative of the growth curve) rather than the growth curve itself. All these references, with the exception of Garre et al. [2023], are primarily concerned with the modeling of a single growth curve. In hierarchical, longitudinal studies, multiple growth curves are generated as a result of monitoring multiple subjects from a group (or multiple groups). In such cases the interest lies in estimating the group-mean-curves, after accounting for the subject specific differences. For these types of problems the dominant approaches have been based on linear mixed-effects, nonlinear mixed-effects, and functional mixed-effects modeling, e.g., see Fine et al. [2019], Blozis and Harring [2016]. In these references, when linear mixed-effects or functional mixed-effects models are used, the group-mean-curves are often linear or polynomial smoothing splines, which, while flexible, do not use the prior monotonicity information of the growth curves. In nonlinear mixed-effects modeling, it is assumed that the subject curves are generated by parametric growth curve models such as Logistic and some or all of the model parameters, like the growth rate, for each subject are assumed to be random perturbations of their mean population values. However, in this case, the subject curves being fully parametric can be a restrictive assumption.
@@
The group-mean-curve, which is expected to follow a growth curve pattern, is called a functional fixed-effect, the deviation-from-mean-curve is called a functional random-effect, regression-effects can either be fixed or random, and the noise is a zero-mean, Gaussian error term. In this article, the group-mean-curve is modeled semiparametrically according to a simple modification of the state space model based approach described in section 3 of Ansley et al. [1993]. Our modification makes their approach computationally feasible in much wider settings. Making their elegant growth curve modeling approach easily accessible to the wider applied research community is one of the goals of this article. All the FME models considered in this article are formulated as linear, Gaussian, state space models, SSMs for short. This provides many additional advantages such as modeling flexibility, computational efficiency, and easy interpretability (see Guo [2002], Selukar [2015]). The analysis for the illustrations in this article is carried out by using the CSSM procedure in the SAS Viya/Econometrics software, see SAS . The data and the SAS code used for these illustrations is available on request.
@@ 2 State Space Modeling Framework
SSMs have long been used to model a variety of sequential data, such as time series, panels of time series, and longitudinal data. Many different forms of SSMs are used in practice. In order to handle the formulation of FME models for the longitudinal data, this article uses an SSM form that is similar to the one described in chap. 6, sec. 4 of Durbin and Koopman [2012]. The notation and additional details are included in the appendix, Appendix A.
@@ 2.1 SSM Formulation of a Parametric Growth Curve
where $\epsilon_{t}$ is a noise sequence of independent, zero-mean, Gaussian variables with variance $\sigma^{2}_{\epsilon}$, and the latent curve, $f(\boldsymbol{\theta},t)=constant+scale\;g(\boldsymbol{\theta},t)$, is a smooth, monotonic function of $t$, which might depend on some parameter vector $\boldsymbol{\theta}$; $constant$ and $scale$ are auxiliary parameters that do not depend on $\boldsymbol{\theta}$. A few polpular choices of $g(\boldsymbol{\theta},t)$ are shown in Table 1; several more examples can be found in Zwietering et al. [1990].
@@
As an illustration, consider the greek_tractors data set considered in Garre et al. [2023]. It contains the number of tractors registered in Greece in a 45 year period between 1961 to 2006. They fit the Logistic curve to $y=\log_{10}(\text{number of tractors})$, by nonlinear least squares. After substituting the parameter estimates, their fitted Logistic curve has the following expression:
@@
Note that, in the SSM formulation only $\phi,\rho$ and $\sigma^{2}_{\epsilon}$ are estimated by the maximum likelihood. The other two parameters, $constant$ and $scale$, are obtained as a result of Kalman smoothing. This is because, Kalman smoothing provides the estimates of the two-dimensional latent state $\boldsymbol{\alpha}_{t}$ at all times and $scale$ is the second element of $\boldsymbol{\alpha}_{t}$. Furthermore, since the first element of $\boldsymbol{\alpha}_{t}$ is $f(t)$, once $scale$, $\phi$, and $\rho$ are known, $constant$ can be derived from the estimate of $f(t)$ at some convenient $t$, say at $t=0$. Finally, Figure 1 shows the estimated Logistic fit (which matches the fit in Garre et al. [2023]) with point-wise 95% confidence band, which is also provided by the Kalman smoother.
@@ 2.2 SSM Formulation of a Semiparametric Growth Curve
The state space formulations in Equation 2 and Equation 5 are based on the treatment of semiparametric growth-curve modeling described in Section 3 of Ansley et al. [1993]. There they describe how to obtain $\mathbf{Q}_{t}$ for different choices of $g(\boldsymbol{\theta},t)$ so that the estimated latent curve, $\mu_{t}$, in Equation 4 satisfies a penalized least-squares criteria. However, except for some simple choices of $g(\boldsymbol{\theta},t)$, obtaining a closed form expression for $\mathbf{Q}_{t}$ is complicated. Instead, in this article we suggest always using the following greatly simplified form of $\mathbf{Q}_{t}$:
@@
  • The transition matrix, $\mathbf{T}_{t}^{t+h}$, used in the formulations in Equation 2 and Equation 5, is exactly the same as in Section 3 of Ansley et al. [1993].
@@
  • For the exponential case, $g(\boldsymbol{\rho},t)=\exp(-\rho\;t)$, a closed form expression for $\mathbf{Q}_{t}$ is described in Ansley et al. [1993] and an equivalent form of this model is discussed in De Jong and Mazzi [2001]. For this exponential case, we compared the growth/decay curves estimated using the exact form of $\mathbf{Q}_{t}$ versus the simplified form in Equation 6 in many real and simulated data cases. In all cases the estimated curves by the exact and approximate methods matched very well. In particular, for the chlorine content data example discussed in both Ansley et al. [1993] and De Jong and Mazzi [2001], the estimated curves were essentially identical.
@@ 2.3 FME Model With Growth Curve As a Functional Fixed-Effect
As a quick illustration of FME model based data analysis, we analyze a small portion of the data used in Kurokawa et al. [2016]. Slightly more elaborate analysis of these data is done in Section 3. Here we consider 12 independent replicate curves generated by monitoring the growth of a bacterium strain in a nutrient rich medium (strain 1, medium LB). The measurements are taken every 30 minutes during a period of 23 hours. Figure 3 shows the growth patterns for replicates 1 and 7. It shows that both the replicates loosely follow the usual sigmoidal shape, however, they do show noticeable differences in their growth patterns.
@@ 3 Illustration: Comparing Bacterial Growth in Different Settings
In Kurokawa et al. [2016], correlation between genome reduction and bacterial growth is studied (a link to access their data is also provided in the article). In their study, E. Coli cultures are grown in three settings of growth mediums and at 36 different levels of gene deletion. For each combination of growth medium and gene deletion level, multiple replicates of E. Coli cultures are grown. Their growth is measured at 30 minute or one hour intervals for different lengths of times. The three growth mediums are M63, MAA, and LB, which indicate poor, supplementary, and rich growth conditions, respectively. The 36 gene deletion levels are labeled as strains 1 to 36, which indicate increasing severity of gene deletion. In our illustration we will analyze the growth patterns for strains 1 and 36, grown at all the three mediums, M63, MAA, and LB (that is, total of 6 strain-medium combinations). Except in the case of (strain 36, medium M63) combination, which has 24 replicates, all other combinations have 12 replicates. Even though our methodology can easily handle different number of replicates for different combinations, we excluded the last 12 replicates for the (strain 36, medium M63) combination. Additionally, two minor adjustments to the data were made for easier viewing of the results:
@@ 4 Conclusion
We have described a flexible SSM-based framework for semiparametric modeling of growth curve data that are generated by hierarchical, longitudinal studies. Our modeling approach is based on a simplification of growth curve modeling approach that is described in Ansley et al. [1993]. We hope our treatment of their approach is more accessible and easier to adopt. The SSM-based framework described here can be easily extended in many directions. For example, the FME model in Equation 7 can contain regression effects, or additional effects that modify the mean curve to account for hormetic effects.
@@ Appendix A SSM Framework and Notation
where the terms $\boldsymbol{\mu}_{t}$, $\boldsymbol{\gamma}_{t},\cdots$ might represent a time-varying mean-level, a seasonal pattern, and so on. Such linear combinations of the state subblocks are called components. When the data from a longitudinal study is assumed to follow an SSM, the data analyis is greately helped by the well-known (diffuse) Kalman filter and (diffuse) Kalman smoother algorithms. Chapters 4, 5, 6, and 7 of Durbin and Koopman [2012] explain how these algorithms provide the following: